Logiweb dictionary of propositional calculus

Up Help

0 0 propositional calculus

1 0 T_{Prop} 2 0 A1 3 0 A2 4 0 A3 5 0 A4 6 0 A5 7 0 MP 8 0 Gen 9 0 Mendelson 1.8 10 0 Mendelson 1.10a 11 0 Mendelson 1.10b 12 0 Mendelson 1.11a

13 0 nani teta saka...:13

14 2 ∗ ⇒ ∗

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-06-02.UTC:18:16:36.700631 = MJD-53523.TAI:18:17:08.700631 = LGT-4624453028700631e-6